OpenStudy (anonymous):

Jackie has two jobs. She earns $14/h in Job A and $16/h in Job B. She would like to earn $560 in a week. Let a represent the number of hours she works in Job A, and let b represent the number of hours she works in Job B. Which function relates the number of hours she works in Job B to the number of the hours she works in Job A? A. b= -8/7a + 40 B.b= -7/8 + 35 C.b= 7/8a - 35 D.b= 8/7a - 40 please help will fan and medal

OpenStudy (anonymous):

@iGreen can you help?

OpenStudy (igreen):

Okay, so we have the equation: 14a + 16b = 560 We have to solve for 'b', so first we subtract 14a to both sides: 16b = -14a + 560 Now divide 16 to both sides: b = -14/16a + 560/16 Divide 560 / 16, what do you get?

OpenStudy (anonymous):

35?

OpenStudy (igreen):

Yep, that gives us: b = -14/16a + 35 Now can you simplify -14/16? Divide both the top and bottom by 2.

OpenStudy (anonymous):

-7/8

OpenStudy (igreen):

Yep, so our final answer is: b = -7/8a + 35

OpenStudy (anonymous):

thank you :) and do you mind checking my answers for three other problems?

OpenStudy (igreen):

No, I don't. Go ahead and post them.

OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (6, 2) and has a slope of 3? y = 3x +16 What is the slope-intercept form of the function that contains the points (5, 6) and (9, 14)? y =2x + -4 What is the slope-intercept form of the function described by this table? x 1 2 3 4 y 2 7 12 17 y = 5x + -3

OpenStudy (igreen):

All are correct except for the first..

OpenStudy (igreen):

We can plug the slope and the point into point-slope form. \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value of the point, \(x_1\) is the x-value of the point, and \(m\) is the slope. \(y - 2 = 3(x - 6)\) Distribute 3 into the parenthesis: \(y - 2 = 3x - 18\) Add 2 to both sides, what's -18 + 2?

OpenStudy (anonymous):

-16

OpenStudy (igreen):

Yes, so our final answer is: \(y = 3x - 16\) or \(y = 3x + -16\)

OpenStudy (anonymous):

thank you